In a gravity free space, a man of mass M standing at a height h above the floor, throws a ball of mass m straight down with velocity u. When the ball reaches the floor, the distance of the man above the floor will be??

Grade:11

1 Answers

DR STRANGE
7 years ago
HI ARIN,
 this is the concept of center of mass
initially the center of mass of them is at height h above the floor, after he he throws the ball the center mass should remain stationary.
let the distance of the man from th floor as ball reaches the floor be h+x (DISTANCE BETWEEN MAN AND THE BALL)
 the distance of center of mass from the ball at floor is h.
mathematically, M\times \frac{h+x}{M+m} = h
solve it for x, x=\frac{hm}{M} 
distance of the man above the floor will be h+x (just add them)
I hope you will get it
plz approve if you are satisfied :)